# MonoGSDF: Exploring Monocular Geometric Cues for Gaussian Splatting-Guided Implicit Surface Reconstruction

### v3

@@ 4.3 Novel View Synthesis
We further compare our method with state-of-the-art novel view synthesis (NVS) methods on the Mip-NeRF 360 dataset [3]. Quantitative results in Table 3 indicate that our approach surpasses GOF [48] across all metrics in indoor scenes, while for outdoor scenes, the performance gap remains relatively small. However, our method significantly improves surface reconstruction quality. This suggests that our enhanced geometry representation leads to higher visual quality. Figure 5 presents a qualitative comparison of the extracted meshes. Consistent with our observations on the Tanks and Temples dataset [17], our method reconstructs smooth, detailed meshes. In contrast, 2DGS [12] yields overly smooth meshes, and GOF [48] produces noisy artifacts on flat surfaces.
4.4 Ablation Study
Figure 6: Ablation Study on Geometry Regularization.
4.4.1 Runtime Analysis
Table 4 presents a comprehensive runtime analysis on the Barn scene, highlighting our method’s efficiency and adaptability. Notably, our framework seamlessly integrates with existing Gaussian rasterizers, as demonstrated by the “Ours (2D)” implementation combining our approach with the 2DGS rasterizer. While the incorporation of SDF guidance enables superior surface reconstruction quality, our method maintains competitive training times comparable to efficient baselines like 2DGS and GOF. Furthermore, our approach achieves substantial acceleration in mesh extraction compared to existing methods.
4.4.2 Ablations on Different Regularizations
Ablations on SDF Fields. Table 5 quantitatively validates the effectiveness of our Gaussian-based implicit surface reconstruction framework compare with GOF. To further demonstrate the generalizability of our approach, we additionally evaluate its performance when integrated with the 2DGS rasterizer denoted as “Ours (2D)”, reporting consistent F1-score improvements across different configurations.
@@ 4.4.2 Ablations on Different Regularizations
Table 4: Runtime. We report the training and meshing time.
Method 2DGS Ours (2D) $\lambda_{D}=0$ $\lambda_{N}=0$ $\lambda_{D}=0$, $\lambda_{N}=0$
F1 Score 0.36 0.40 0.52 0.51 0.50
Method GOF GOF+Geo $\lambda_{D}=0.5$ $\lambda_{N}=0.5$ Ours
F1 Score 0.51 0.52 0.45 0.47 0.56
Table 5: Quantitative Ablation Study. We report F1 score of scene Barn with and without proposed regularization. We denote our method with default 3D Gaussian rasterizer [48] as “Ours” and with 2D Gaussian rasterizer [12] as “Ours (2D)”. “GOF+Geo” represents GOF with geometry cues only.
Ablations on Geometry Regularization. The ablation studies in Table 5 demonstrate the impact of different geometry regularization terms. Since the pseudo normal map is derived from the pseudo depth map, they exhibit identical performance gaps. We further evaluate GOF enhanced with our geometry cues, denoted as “GOF+Geo”, which shows modest improvements (0.52). Due to the scale ambiguity problem, we set small values for $\lambda_{D},\lambda_{N}$ specifically (0.05 and 0.1 for Tanks and Temples, 0.01 and 0.01 for DTU and MipNeRF 360). We demonstrate that increasing these weights to 0.5 actually degrades the results by introducing adverse effects from geometric cues. We also demonstrate the error distribution after alignment in supp. mat., showcasing the limitation of geometry cues. The most significant enhancement is achieved through our complete pipeline (8% improvement), highlighting the synergistic effect of our integrated approach. Figure 6 demonstrates the effectiveness of our geometry regularization in handling transparent and flat regions, significantly improving reconstruction quality through enhanced surface consistency.
@@ 5 Conclusion
We introduce MonoGSDF, a Gaussian-based implicit surface reconstruction framework. Our SDF-to-Opacity function bridges implicit and explicit representations, while Gaussian-guided normalization ensures stable optimization across unbounded 3D scenes. A multi-resolution training strategy, enhanced by the geometric cues, progressively refines reconstruction quality. Our method achieves state-of-the-art accuracy and efficiency, effectively handling complex cases such as flat and transparent surfaces where existing approaches often struggle.
Supplementary Material
@@ A.1 Hyperparameters
Gaussian Splatting. For hyperparameters used in the Gaussian rasterization, we follow previsous works [48, 12, 16]. We train our Gaussian model and SDF network jointly with 30000 iterations. We set the initial learning rate for Gaussians’ position as 0.00016 and the final initial learning rate as 0.0000016. And we set learning rates for scales and rotation as 0.005, 0.001, respectively. We start the Gaussian densification after 500 iterations and until 15000 iterations. We densify Gaussians every 100 iterations and the gradient threshold for densification is 0.0002. We start the distortion loss $\mathcal{L}_{d}$ and the depth-normal loss $\mathcal{L}_{n}$ after 3000 iterations and 7000 iterations, respectively. And we set the wavelet level as 3 and gradually increase the resolution until 10000 iteration. After that, we use the full resolution for training.
@@ Appendix B Distortion and Depth-Normal Loss
We apply a distortion loss [12] and depth-normal loss [48] as discussed in Section 3.2. The distortion loss concentrates the weight distribution along the rays by minimizing the distance between the ray-splat intersections:
@@
Appendix C Additional Experimental Results
C.1 Results on DTU and Mip-NeRF 360
In Table 7, we present the results reported in the Neuralangelo publication [21] on the DTU dataset. It is important to note that the quantitative results from the Neuralangelo paper have been flagged by other studies and discussions on the official GitHub repository as non-reproducible [^2] Refer to this GitHub issue for details..
To provide a more comprehensive evaluation, we include additional qualitative comparisons on the DTU dataset [13] in Figure 13 and further demonstrate results on the Mip-NeRF 360 dataset [3] in Figure 14. Our method excels in handling reflective surfaces and produces reconstructions that are not only smoother but also exhibit finer details, highlighting its capability for more accurate and visually appealing scene representations.
Appendix D More Ablation
@@
D.1 Ablation on Different Rasterizers
In Figure 7, we present reconstruction results using different Gaussian rasterizers, showcasing the versatility and adaptability of our method. These results demonstrate that our approach can be seamlessly integrated into existing Gaussian methods, no matter it’s 2D Gaussians [12] or 3D Gaussians [48], enhancing their performance without requiring significant modifications. Furthermore, our method consistently delivers high-quality reconstructions with fine details, highlighting its effectiveness in capturing intricate scene structures such as transparent and reflection areas.
D.2 Ablations on Normalization Function.
We acknowledge that non-uniform transformation like MipNeRF’s [3] non-linear normalization (mapping $\pm\infty$ to (2,2))(-2,2)) can cause distortions and conflicts with tinycudann’s requirements of having inputs in $(0,1)$. In contrast, our method applies a near-linear transformation for Gaussians within the initial bounding box and a compression for outliers, minimizing distortion effects. As shown in Table 6, our approach (“Ours”) outperforms MipNeRF’s normalization function (“w/ MipNeRF NF”) which is also scaled to $(0,1)$.
D.3 Ablations on SDF to Opacity Function.
We compare results using 3DGSR’s [24] SDF-to-Opacity function (denoted as “w/ 3DGSR SDF2O” in Table 6) with our proposed function (Eq. 5). Our method achieves a higher f-score, offering a simpler, more efficient solution and the ability to reach full opacity of 1, unlike 3DGSR’s capped opacity of 0.25. While 3DGSR targets object-level reconstruction, this limitation may impede its effectiveness in unbounded outdoor scenes.
Method w/ MipNeRF NF w/ 3DGSR SDF2O Ours
F-score 0.52 0.24 0.56
Table 6: Ablation Study on Normalization and SDF to Opacity. We show the F-score of reconstructed mesh on Barn.
D.4 Ablations on Geometry Cues.
Our analysis reveals significant artifacts in Depth Anything v2’s estimated depth maps due to inherent scale uncertainty, despite global alignment attempts. As shown in Figure 8, we compare our alignment method “Ours” with DN Splatter’s approach [34] “DN Align”, demonstrating our method’s superior ability to reduce depth errors between rendered and aligned pseudo depth. Interestingly, despite the inaccuracies in DN Splatter’s aligned pseudo depth, the reconstruction F1 score remains comparable to our method. This finding suggests that while these limitations underscore the unreliability of pseudo depth maps for precise reconstruction, necessitating our model’s strong ability to compensate for misalignments, the primary improvement stems from our novel Gaussian-SDF collaboration pipeline rather than geometric cues alone.
Figure 8: Ablations on Geometry Cues. We show the ground truth color image, rendered depth, aligned pseudo depth and depth error (in log) on Barn. DN Align represents our pipeline with DN Splatter’s [34] alignment.
D.5 Ablations on Pruning Strategy.
@@ D.1 Ablation on Different Rasterizers
D.6 Ablations on Multi-Resolution Regularization.
@@
D.7 Ablations on Different Meshing Methods.
Figure 11 illustrates the limitations of depth fusion-based mesh extraction in unbounded scenes, revealing significant distortion caused by resolution constraints and MipNeRF’s [3] normalization artifacts. Comparative results in Figure 12 demonstrate our method’s advantages across different meshing techniques, showing that our Gaussian-guided surface reconstruction achieves both superior efficiency and enhanced detail preservation compared to existing approaches.
@@
Appendix E Limitation Discussion
@@
N-angelo [21] 0.49 1.05 0.95 0.38 1.22 1.10 2.16 1.68 1.78 0.93 0.44 1.46 0.41 1.13 0.97 1.07
N-angelo* [21] 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61
@@
Table 7: Quantitative Comparison on the DTU Dataset [13]. We show the Chamfer distance. For Neuralangelo [21], we report the results from UniSDF [35] reproduction as N-angelo, and the results from Neuralangelo publication as N-anglo*, which is not reproducible.
Figure 13: Additional Results on DTU [13].
Figure 14: Additional Results on Mip-NeRF 360 [3].

### v4

@@ 4.3 Novel View Synthesis
We further compare our method with state-of-the-art novel view synthesis (NVS) techniques on the Mip-NeRF 360 dataset [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman]. The quantitative results presented in Table 3 demonstrate that our approach achieves remarkable rendering performance. Our method effectively aligns the Gaussians with the surface, leading to better surface reconstruction without compromising rendering quality.
@@
Table 4: Runtime. We report the training and meshing time. Compared to GOF [Yu et al.(2024c)Yu, Sattler, and Geiger], we avoid ray tracing for SDF evaluation, enabling fast meshing.
Method 2DGS Ours (2D) $\lambda_{D}=0$ $\lambda_{N}=0$ $\lambda_{D}=0$, $\lambda_{N}=0$ w/ MipNeRF NF
F1 Score 0.36 0.40 0.52 0.51 0.50 0.52
Method GOF GOF+Geo $\lambda_{D}=0.5$ $\lambda_{N}=0.5$ Ours w/ 3DGSR SDF2O
F1 Score 0.51 0.52 0.45 0.47 0.56 0.24
Table 5: Quantitative Ablation Study. We report Barn scene F1-scores with and without different regularization.
4.4 Runtime and Ablation Study
Figure 5: Ablation Study on Geometry Regularization. We show the reconstructed meshes with and without the geometry regularization.
Runtime Analysis.
Table 4 reports the runtime analysis on the Barn scene. Our method integrates seamlessly with different Gaussian rasterizers. “Ours (2D)” refers our method to using the 2DGS rasterizer. While SDF guidance enhances surface reconstruction, our approach achieves competitive training times and notably faster mesh extraction than baselines such as 2DGS and GOF, since our SDF network directly outputs values for any query point, whereas GOF relies on ray tracing and 2DGS requires TSDF fusion.
Ablations on Geometry Regularization.
Table 5 presents ablation studies on the impact of different geometry regularization terms, showing similar performance gaps for pseudo normal and depth maps. Evaluating GOF with geometry cues only ("GOF+Geo") shows modest improvements (0.52). To highlight that geometry cues alone aren’t the main factor, we increase $\lambda_{D}$ and $\lambda_{N}$ to 0.5, leading to performance degradation due to their adverse effects. Our complete pipeline achieves an 8% improvement, demonstrating the synergistic effect of our approach. Figure 5 illustrates the effectiveness of our geometry regularization, especially in improving surface consistency and reconstruction quality in transparent and flat regions.
Ablations on Normalization Function.
MipNeRF 360’s non-linear normalization [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman] introduces distortions and conflicts with tinycudann’s $(0,1)$ input requirement. In contrast, our near-linear normalization within the bounding box compresses outliers with minimal distortion. For fairness, we scale MipNeRF’s normalization output to $(0,1)$ and compare with ours in Table 5, where our method ("Ours") achieves superior performance.
Ablations on SDF to Opacity Function.
We compare 3DGSR’s SDF-to-Opacity function [Lyu et al.(2024)Lyu, Sun, Huang, Wu, Yang, Chen, Pang, and Qi] ("w/ 3DGSR SDF2O") with our method in Table 5. Our approach achieves a higher F1-score, providing a simpler, more efficient solution with full opacity, while 3DGSR’s opacity is capped at 0.25, limiting its effectiveness in unbounded outdoor scenes.
Ablations on SDF Fields.
Table 5 shows that our Gaussian-based implicit surface reconstruction outperforms GOF, and when integrated with the 2DGS rasterizer (“Ours (2D)”), it consistently achieves higher F1-scores across settings.
@@ 5 Conclusion
We propose MonoGSDF, a Gaussian-based implicit surface reconstruction framework that bridges implicit and explicit representations via an SDF-to-opacity mapping. Gaussian-guided normalization stabilizes optimization in unbounded scenes, and multi-resolution training with geometric cues enables progressive and accurate reconstruction. Our method achieves SOTA performance, especially on flat and transparent surfaces where prior methods fail.
@@ A.1 Hyperparameters
Gaussian Splatting. For hyperparameters used in the Gaussian rasterization, we follow previsous works [Yu et al.(2024c)Yu, Sattler, and Geiger, Huang et al.(2024)Huang, Yu, Chen, Geiger, and Gao, Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis]. We train our Gaussian model and SDF network jointly with 30000 iterations. We set the initial learning rate for Gaussians’ position as 0.00016 and the final initial learning rate as 0.0000016. And we set learning rates for scales and rotation as 0.005, 0.001, respectively. We start the Gaussian densification after 500 iterations and until 15000 iterations. We densify Gaussians every 100 iterations and the gradient threshold for densification is 0.0002. We start the distortion loss $\mathcal{L}_{d}$ and the depth-normal loss $\mathcal{L}_{n}$ after 3000 iterations and 7000 iterations, respectively. And we set the wavelet level as 3 and gradually increase the resolution until 10000 iteration. After that, we use the full resolution for training.
@@ Appendix B Distortion and Depth-Normal Loss
We apply a distortion loss [Huang et al.(2024)Huang, Yu, Chen, Geiger, and Gao] and depth-normal loss [Yu et al.(2024c)Yu, Sattler, and Geiger] as discussed in Section 3.2. The distortion loss concentrates the weight distribution along the rays by minimizing the distance between the ray-splat intersections:
@@
Appendix C Mesh Extraction.
Previous Marching Cubes-based methods [Yu et al.(2024a)Yu, Lu, Xu, Jiang, Xiangli, and Dai, Lyu et al.(2024)Lyu, Sun, Huang, Wu, Yang, Chen, Pang, and Qi] require dense querying across both occupied and free space, resulting in inefficiencies and limited resolution—especially in unbounded scenes with coordinate distortions. In contrast, we adopt Marching Tetrahedra [Kulhanek and Sattler(2023)] and leverage spatial cues from 3D Gaussians to restrict the search to regions near the surface, significantly reducing unnecessary computation while improving surface alignment and reconstruction efficiency. Unlike GOF [Yu et al.(2024c)Yu, Sattler, and Geiger], which performs costly ray tracing to query opacity values for each point, our method directly outputs SDF values, accelerating meshing from 30 minutes to just 5 minutes.
Appendix D Additional Experimental Results
D.1 Results on DTU and Mip-NeRF 360
In Table 6, we present the results reported in the Neuralangelo publication [Li et al.(2023)Li, Müller, Evans, Taylor, Unberath, Liu, and Lin] on the DTU dataset. It is important to note that the quantitative results from the Neuralangelo paper have been flagged by other studies and discussions on the official GitHub repository as non-reproducible [^2] Refer to this GitHub issue for details..
To provide a more comprehensive evaluation, we include additional qualitative comparisons on the DTU dataset [Jensen et al.(2014)Jensen, Dahl, Vogiatzis, Tola, and Aanæs] in Figure 14 and further demonstrate results on the Mip-NeRF 360 dataset [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman] in Figure 6 and 7. Our method excels in handling reflective surfaces and produces reconstructions that are not only smoother but also exhibit finer details, highlighting its capability for more accurate and visually appealing scene representations.
Figure 6: Surface Reconstruction on the Mip-NeRF 360 Dataset [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman]. We show the rendered normal maps from reconstructed meshes.
Figure 7: Additional Results on Mip-NeRF 360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman].
Appendix E More Ablation
@@
E.1 Ablation on Different Rasterizers
In Figure 8, we present reconstruction results using different Gaussian rasterizers, showcasing the versatility and adaptability of our method. These results demonstrate that our approach can be seamlessly integrated into existing Gaussian methods, no matter it’s 2D Gaussians [Huang et al.(2024)Huang, Yu, Chen, Geiger, and Gao] or 3D Gaussians [Yu et al.(2024c)Yu, Sattler, and Geiger], enhancing their performance without requiring significant modifications. Furthermore, our method consistently delivers high-quality reconstructions with fine details, highlighting its effectiveness in capturing intricate scene structures such as transparent and reflection areas.
E.2 Ablations on Geometry Cues.
Our analysis reveals significant artifacts in Depth Anything v2’s estimated depth maps due to inherent scale uncertainty, despite global alignment attempts. As shown in Figure 9, we compare our alignment method “Ours" with DN Splatter’s approach [Turkulainen et al.(2024)Turkulainen, Ren, Melekhov, Seiskari, Rahtu, and Kannala] “DN Align", demonstrating our method’s superior ability to reduce depth errors between rendered and aligned pseudo depth. Interestingly, despite the inaccuracies in DN Splatter’s aligned pseudo depth, the reconstruction F1 score remains comparable to our method. This finding suggests that while these limitations underscore the unreliability of pseudo depth maps for precise reconstruction, necessitating our model’s strong ability to compensate for misalignments, the primary improvement stems from our novel Gaussian-SDF collaboration pipeline rather than geometric cues alone.
Figure 9: Ablations on Geometry Cues. We show the ground truth color image, rendered depth, aligned pseudo depth and depth error (in log) on Barn. DN Align represents our pipeline with DN Splatter’s [Turkulainen et al.(2024)Turkulainen, Ren, Melekhov, Seiskari, Rahtu, and Kannala] alignment.
E.3 Ablations on Pruning Strategy.
@@ E.1 Ablation on Different Rasterizers
E.4 Ablations on Multi-Resolution Regularization.
@@
E.5 Ablations on Different Meshing Methods.
Figure 12 illustrates the limitations of depth fusion-based mesh extraction in unbounded scenes, revealing significant distortion caused by resolution constraints and MipNeRF’s [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman] normalization artifacts. Comparative results in Figure 13 demonstrate our method’s advantages across different meshing techniques, showing that our Gaussian-guided surface reconstruction achieves both superior efficiency and enhanced detail preservation compared to existing approaches.
@@
Appendix F Limitation Discussion
@@
N-angelo [Li et al.(2023)Li, Müller, Evans, Taylor, Unberath, Liu, and Lin] 0.49 1.05 0.95 0.38 1.22 1.10 2.16 1.68 1.78 0.93 0.44 1.46 0.41 1.13 0.97 1.07
N-angelo* [Li et al.(2023)Li, Müller, Evans, Taylor, Unberath, Liu, and Lin] 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61
@@
Table 6: Quantitative Comparison on the DTU Dataset [Jensen et al.(2014)Jensen, Dahl, Vogiatzis, Tola, and Aanæs]. We show the Chamfer distance. For Neuralangelo [Li et al.(2023)Li, Müller, Evans, Taylor, Unberath, Liu, and Lin], we report the results from UniSDF [Wang et al.(2023)Wang, Rakotosaona, Niemeyer, Szeliski, Pollefeys, and Tombari] reproduction as N-angelo, and the results from Neuralangelo publication as N-anglo*, which is not reproducible.
Figure 14: Additional Results on DTU [Jensen et al.(2014)Jensen, Dahl, Vogiatzis, Tola, and Aanæs].